# Synthetic Data Generation for Culturally Nuanced Commonsense Reasoning in Low-Resource Languages

### v1

@@ Appendix H Agreement Score
Fluency 0.74 0.76
Coherency 0.78 0.71
Correctness 0.73 0.50
Cultural Relevance 0.23 0.49
Table 7: Agreement Scores for Javanese and Sundanese
Appendix I Classification Test Performance on Javanese and Sundanese
Tables 8 and 9 present test performance on Sundanese and Javanese across different training datasets. XLM-R and BiLSTM outperform the n-gram and FastText baselines, with LLM-generated datasets achieving comparable or better results than machine translation. Notably, All $LLMs+Filtered$ training sets yield the best performance, especially on human-written test sets, highlighting the benefit of curated synthetic data.
Appendix J Workers Guidelines
Rubric for Story Evaluation
  • Fluency: Each sentence should be grammatically correct and fluent.
    • Score Range: 0 to 5
    • 5: All sentences are grammatically correct and very fluent.
@@ Appendix I Classification Test Performance on Javanese and Sundanese
  • Coherency: The story should be coherent, with all sentences logically connected.
    • Score Range: 0 to 5
@@
  • Correctness: The correct story closure should be valid, while the incorrect closure should clearly be wrong.
    • Binary Score:
@@
  • Cultural Relevance: The story should reflect appropriate cultural norms, values, or symbols relevant to the language.
    • Binary Score:
    • 1: The story contains relevant cultural norms/values/symbols for the corresponding language.
@@
Appendix K Topic Taxonomy and Story-Writing Guidelines
Topic Taxonomy
@@
As part of the Javanese and Sundanese Cloze Project, a total of 300 stories must be created. These stories are evenly distributed across 12 predefined topic categories, with 25 stories per topic. Each story must reflect traditional Javanese and Sundanese values and customs, with attention to detail and coherence in the narrative. The topic categories and their subcategories are as follows:
  • Food: Breakfast, Lunch, Dinner, Snacks, Food souvenirs, Traditional foods and beverages, Eating habits, Cutlery, Cooking ware, Fruits.
  • Wedding: Traditions before marriage, Traditions during marriage, Traditions after marriage, Men’s wedding clothes, Women’s wedding clothes, Invited guests, Wedding location, Foods at a wedding, Gifts brought to a wedding.
  • Family Relationship: Relationships within the immediate family, Relationships in the extended family, Relations with society or neighbors, Clan or descendant system.
  • Pregnancy and Kids: Traditions during pregnancy, Traditions after birth, Caring for a newborn baby, Caring for toddlers, Caring for children, Caring for teenagers, Parent-child interactions as adults.
  • Death: When death occurs, The process of caring for a corpse, Traditions after burial, Mourners’ clothes, Inheritance matters.
  • Religious Holiday: Traditions before religious holidays, Traditions during religious holidays, Traditions after religious holidays.
  • Agriculture: What to plant, Traditions during planting, Harvest.
  • Fisheries and Trade: Traditions of taking care of livestock or fish, Buying and selling traditions.
  • Art: Musical instruments, Folk songs, Traditional dances, Use of art at specific events, Poetry or similar literature.
  • Traditional Games: Types of games, Locations where games are played.
  • Daily Activities: Morning activities, Afternoon activities, Evening activities, Leisure activities, Household tasks, Transportation.
  • Socio-religious Aspects of Life: Regular religious activities, Mystical elements, Traditional ceremonies, Lifestyle, Self-care, Traditional medicine, Traditional sayings.
Story-Writing Guidelines
Each story must consist of 4 sentences and two endings: one correct and one incorrect. The following example demonstrates how stories should be written (Table 10):
Ensure that all stories adhere to the topics and categories outlined in the taxonomy and reflect traditional values and cultural relevance.
Figure 10: Prompt template instructing LLM to generate a new example for Javanese and Sundanese, given a set of in-context examples.
Train Set Machine Translated - Test Human-written
n-gram FastText BiLSTM XLM-R n-gram FastText BiLSTM XLM-R
ID 54.6 63.0 51.4 ± 1.1 58.4 ± 0.6 55.7 62.5 49.3 ± 0.2 68.0 ± 0.6
MT 59.3 ± 0.2 67.2 ± 0.2 57.5 ± 0.1 74.3 ± 0.1
GPT-4o 63.5 ± 0.1 67.2 ± 0.1 62.8 ± 0.1 76.6 ± 0.1
Claude 63.5 ± 0.1 68.0 ± 0.1 63.7 ± 0.1 76.9 ± 0.1
Cohere 57.0 ± 0.1 62.6 ± 0.1 54.3 ± 0.1 70.5 ± 0.2
Llama3.1 60.4 ± 0.2 67.2 ± 0.1 58.1 ± 0.1 73.7 ± 0.1
Mixtral 57.4 ± 0.1 54.4 ± 0.1 59.2 ± 0.2 54.7 ± 0.1
Gemma2 59.4 ± 0.1 55.6 ± 0.1 61.2 ± 0.1 54.3 ± 0.1
All LLMs 63.4 ± 0.1 68.3 ± 0.1 63.1 ± 0.2 75.3 ± 0.1
All LLMS + Filtering 62.0 ± 0.1 67.2 ± 0.1 61.7 ± 0.1 71.5 ± 0.1
Table 8: Test Performance on Sundanese across Different Models and Training Data
Train Set Machine Translated - Test Human-written
n-gram FastText BiLSTM XLM-R n-gram FastText BiLSTM XLM-R
ID 63.5 61.0 49.7 ± 0.1 63.2 ± 0.7 72.4 62.9 40.6 ± 0.1 61.8 ± 0.1
MT 60.4 ± 0.1 70.5 ± 0.3 55.6 ± 0.2 65.7 ± 0.2
GPT-4o 56.7 ± 0.3 63.6 ± 0.1 71.0 ± 0.2 83.7 ± 0.1
Claude 58.9 ± 0.1 63.8 ± 0.1 73.4 ± 0.1 79.5 ± 0.1
Cohere 52.8 ± 0.2 64.2 ± 0.1 64.2 ± 0.1 77.2 ± 0.1
Llama3.1 58.5 ± 0.2 65.3 ± 0.1 63.5 ± 0.1 81.9 ± 0.1
Mixtral 58.1 ± 0.1 56.2 ± 0.1 66.2 ± 0.1 54.5 ± 0.1
Gemma2 52.4 ± 0.2 56.6 ± 0.1 48.3 ± 0.2 59.9 ± 0.1
All LLMs 60.4 ± 0.3 66.0 ± 0.1 73.4 ± 0.1 80.6 ± 0.1
All LLMs + Filtering 53.9 ± 0.1 63.6 ± 0.1 73.0 ± 0.1 79.5 ± 0.1
Table 9: Test Performance on Javanese across Different Models and Training Data
Topic Category Javanese Translated
Javanese
Food Traditional foods and beverages Premise: Laras seneng banget karo wedang uwuh, minuman tradisional saka Yogyakarta. Saben sore, dheweke mesthi ngombe wedang uwuh kanggo njaga kesehatan lan semangat. Wedang uwuh digawe saka campuran rempah-rempah alami sing wangi. Laras percaya yen wedang uwuh bisa ngusir rasa kesel lan lemes. Correct: Saben sore, dheweke seneng bareng-bareng karo kanca-kanca ngombe wedang uwuh. Incorrect: Dheweke ngindhari ngombe wedang uwuh saben sore. Premise: Laras is really happy with wedang uwuh, a traditional drink from Yogyakarta. Every evening, she must always drinks wedang uwuh to maintain her health and vitality. Wedang uwuh is made from a mixture of fragrant natural spices. Laras believes that wedang uwuh can drive away fatigue and weakness. Correct: Every evening, she enjoys drinking wedang uwuh with her friends. Incorrect: She avoids drinking wedang uwuh every evening.
Sundanese
Wedding Traditions when getting married Premise: Kulawargi bapak Syakur nembe pisan hajat nikahan. Anjeunna parantos nikahkeun putrana anu bungsu, Anton. Acara sawéranna heboh pisan, aya rupi-rupi doorprize. Abi emut aya hiji ibu-ibu nu kéngéng rice cooker. Correct: Aya ibu-ibu nu kéngéng doorprize tina sawéran hajatan pak Syakur. Incorrect: Pak Syakur masihan rice cooker kumargi ibu-ibu eta orang Premise: My family recently had a wedding. He has married his youngest son, Anton. The event was very exciting, there were various door prizes. I remember there was a mother who received a rice cooker. Correct: There were mothers who received door prizes from Mr. Syakur’s wedding. Incorrect: Mr. Syakur gave the rice cooker because the mothers were improtant people.
Table 10: Example of a story with correct and incorrect endings.

### v2

@@ Appendix C Agreement Score
Fluency 0.73 0.76
Coherency 0.77 0.71
Correctness 0.74 0.50
Cultural Relevance 0.75 0.49
Appendix D Workers Scoring Guidelines
  • Fluency (0–5): Each sentence should be grammatically correct and fluent.
    • 5: All sentences are grammatically correct and fluent.
@@ Appendix D Workers Scoring Guidelines
  • Coherency (0–5): The story should be coherent, with all sentences logically connected.
@@
  • Correctness (Binary): The correct story closure should be valid, while the incorrect closure should clearly be wrong.
@@
  • Cultural Relevance (Binary): The story should reflect appropriate cultural norms, values, or symbols relevant to the language.
    • 1: The story contains relevant cultural norms, values, symbols for the corresponding language.
@@
Appendix E Topic and Story-Writing Guidelines
@@
We create a total of 300 native-authored stories as part of the Javanese and Sundanese Cloze Project. These stories are evenly distributed across 12 predefined topic categories, with 25 stories per topic. Each story must reflect traditional Javanese and Sundanese values and customs, with attention to detail and coherence in the narrative. The topic categories and their subcategories follows IndoCulture Koto et al. (2024). Each story must consist of 4 sentences and two endings: one correct and one incorrect. Ensure that all stories adhere to the topics and categories outlined in the taxonomy and reflect traditional values and cultural relevance.
Appendix F Machine Translation Evaluation
Javanese Sundanese
Google Translate 0.11 / 0.41 / 45.2 0.09 / 0.35 / 42.8
NLLB-200-3.3B 0.18 / 0.30 / 53.1 0.12 / 0.33 / 43.8
Table 9: BLEU F1 / METEOR / ChrF comparison of Google Translate and NLLB Dense Transformers Translation Model Costa-Jussà et al. (2022) for Indonesian to Javanese / Sundanese translation on NusaX dataset Winata et al. (2023)